OpenStudy (anonymous):

Show all steps used to solve the problem. One point for steps and one point for solution. 17-x over 4 =-10

OpenStudy (jackellyn):

17-x/4 =-10 17 -x /4 *4 = -10 *4 17-x = -40 -x = -40 -17 x =57

OpenStudy (anonymous):

Whoo!! Thanks girlie! :]

OpenStudy (jackellyn):

np